### v1

# The 4th Dimension for Scaling Model Size

Abstract
Scaling the size of large language models typically involves 3 dimensions: depth, width, and the number of parameters. In this work, we explore the 4th dimension, virtual logical depth (VLD), which allows increasing the effective algorithmic depth without changing the overall parameter count by reusing parameters within the model. Although parameter reuse itself is not new, its intriguing potential and characteristics in model scaling have not been thoroughly studied. We carefully design controlled experiments and have the following key discoveries on VLD scaling:
  1. 1.
    VLD scaling forces the knowledge capacity of the model to stay almost constant, though with some non-significant variations.
  2. 2.
    VLD scaling enables the reasoning capability to be significantly improved, if the scaling method is properly implemented.
  3. 3.
    The number of parameters is proportional to knowledge capacity, but not reasoning capability. Under certain conditions, it is not necessary to increase parameter count to improve reasoning.
  4. 4.
    The above observations hold for various model configurations and are likely to be generally true under the scope of our experiments.
These findings not only provide useful insights on the future model scaling strategies, but also introduces a even deeper question: Do we really need a lot of parameters to get a really intelligent model? We believe there are many unknown dynamics inside model scaling that needs exploration, so we fully open-source our models and code (https://vldscaling.ngrok.io) to ensure reproducibility and facilitate future explorations.
Figure 1: Comparison between classical model size scaling vs. VLD scaling. Bubble size proportional to total effective depth (base layer depth + virtual logical depth). Blue bubbles represent standard models without VLD scaling but following the classical model size scaling, while green bubbles show models with VLD scaling applied, demonstrating near-vertical scaling paths. VLD scaling significantly enhances reasoning capability (y-axis) while keeping the knowledge capacity (x-axis) almost constant without significant variations.
1 Introduction
Scaling up large language models (LLMs) is widely recognized as a key factor in enhancing their general capabilities (Kaplan et al., 2020; Hoffmann et al., 2022; Brown et al., 2020; Wei et al., 2022). The general concept of model scaling several directions, including model size scaling, data scaling, compute scaling (Wei et al., 2022), and the recent inference scaling (Snell et al., 2024; Jaech et al., 2024). The model size scaling direction typically occurs along three dimensions: depth, width, and the total number of parameters  [^1] Since depth and width directly influence the number of parameters, these three dimensions are not entirely independent. However, the total parameter count can indeed be adjusted independently by modifying the inner dimension of the MLP layer within a transformer block.. Contemporary LLMs often contain an extremely large number of parameters, sometimes exceeding 100 billion (Wei et al., 2022). While such models can store vast amounts of knowledge, this does not necessarily equate to strong reasoning capabilities. In contrast, the human brain exhibits the opposite characteristics. Human memory is relatively limited—hence the invention of tools like pen and paper to compensate for this weakness—yet humans excel at reasoning (Chollet, 2019; Xu & Poo, 2023; Wang & Li, 2024). This cognitive ability is largely supported by a specialized reasoning core located in the prefrontal cortex (Raichle, 2009).
This fundamental difference in the strengths and weaknesses of language models and the human brain raises critical questions about the efficacy of the current scaling paradigm in achieving super-intelligence. Ideally, a truly intelligent system would consist of a powerful reasoning core complemented by a selective yet essential memory, with additional domain-specific knowledge retrieved externally as needed (Kumaran et al., 2016).
This insight motivates us to explore new scaling dimensions that intentionally limits the knowledge capacity to force it towards improving reasoning capabilities. We find that simple weight-sharing within the standard transformer architecture performs surprisingly well. This introduces the core concept of this manuscript: virtual logical depth (VLD). VLD is achieved by repeating transformer layers while sharing weights across these repetitions. Although not entirely new, the intriguing properties of VLD in model scaling is not thoroughly explored. Previous studies (Lan et al., 2020; Dehghani et al., 2019; Bai et al., 2019; Hay & Wolf, 2024; Bhojanapalli et al., 2021; Reid et al., 2021; Takase & Kiyono, 2021; Liu et al., 2023) have discussed parameter reuse or layer repetition in different contexts, but their primary focus was not on understanding its dynamics in scaling. The looped transformers (Saunshi et al., 2025) is a pioneering work that studies the reasoning capability among the related work, but we focus more on the model scaling strategy and provide a set of more comprehensive experiments and new conclusions on the VLD scaling law.
Through carefully controlled experiments, we investigated the characteristics of VLD scaling, specifically measuring reasoning capabilities and knowledge capacity under various configurations. We discovered that VLD scaling shows very different properties from classical model size scaling. In brief, VLD scaling forces the knowledge capacity of the model to stay almost constant (with non-significant variations), while significantly improving the reasoning capability. We illustrate a part of our main findings in Figure 1, which compares the two scaling strategies. These findings hold for various model and experiment configurations, and are likely to be universal under our scope of experiments. Under reasonable conditions, we don’t have to increase the number of parameters in order to increase the inherent reasoning capability of the model.
Beyond the scope of this work, we would like to raise a even deeper question: Do we really need a lot of parameters to achieve super-intelligence? Of course this first depends on how we define super-intelligence. Humans have poor memory but good reasoning capability and agency capability (e.g., subjective initiatives). Under limited human brain capacity, why did the natural evolution chose to allocate more capacity on reasoning capability and agency capability, but only some essential capacity for memory? It is not impossible that under a small but essential amount of parameters and memory capacity, we will be able to push the reasoning capability and agency capability to the extreme and achieve super-intelligence in the far future. We would like to move a small step towards this ultimate goal, and open-source all source code of our current work to facilitate future research.
Figure 2: Different patterns of parameter reuse to increase VLD while keeping the total number of parameters constant. (a) an example of a standard transformer with 3 layers. (b) sequentially repeat neighboring layers. (c) cycle-repeat the layers. (d) repeat layers in an inverse-cycle order. Under the same pattern, two layers with the same color share the same parameters. In all cases, the actual number of parameters do not change.
2 Virtual Logical Depth
VLD is the effective algorithmic depth of the model minus the number of base layers. See Figure 2 for a brief illustration. In this example, there are 3 base layers and 3 VLD layers. We reuse the base layer parameters in different patterns (Takase & Kiyono, 2021) and produce different types of VLD layers. While many previous work on parameter reuse (Reid et al., 2021; Liu et al., 2023) attempted to modify the standard transformer architecture, our focus is not a new architecture but a new scaling dimension under a general architecture. Therefore, we keep the standard transformer architecture and only repeat its layers while sharing their weights. We consider three patterns of parameter reuse: sequence, cycle and inverse cycle, as illustrated in Figure 2 (b-d). In the sequence mode, the neighboring layers share the same parameter. In the cycle mode, a number of layers are grouped into a large block and then the block is repeated. In the inverse cycle mode, a number of layers are grouped into a large block but the block is reversed when it is repeated. We define the core question that we want to explore as follows:
Problem Statement.
Given the VLD scaling and standard model size scaling, find the difference of their influence on knowledge capacity and reasoning capability of the model.
The above question requires us to develop reasonable approaches to measure the knowledge capacity and reasoning capability. Our high-level guideline is that such kind of measurement should be done with rigorously controlled experiments like in Allen-Zhu & Li (2024). We got inspiration from Allen-Zhu & Li (2024) and develop the measurement approaches as explained in Section 2.1 and Section 2.2.
2.1 Measurement of Knowledge Capacity
Knowledge capacity can estimated by measuring the maximum information entropy (Gray, 2011; Volkenstein, 2009) that a LLM absorbs. For a discrete variable $X$ that has $n$ possible values $x_{1},x_{2},\cdots,x_{n}$, the expected information entropy (in bits) for a realization of $X$ is:
$H(X)=-\sum_{i=1}^{n}p(x_{i})\mathrm{log_{2}}p(x_{i})$
Given this, we follow these steps to measure the knowledge capacity:
Step 1: Construct a Random Number Dataset.
We construct a random number dataset, which is a sequence of $k$ numbers, where each number is a realization of $X$, uniformly drawn from the $n$ possible values. The information entropy bits of this dataset is:
$H_{1}=-k\sum_{i=1}^{n}\frac{1}{n}\mathrm{log_{2}}\frac{1}{n}=k\mathrm{log_{2}}n$
Step 2: Train a LLM to Remember the Random Number Sequence.
Now assuming we train a LLM to fit the dataset. Each random number corresponds to exactly one token. The softmax output of the LLM is a length-$n$ vector that predicts the probability of the next token. If the LLM is not trained at all, the softmax vector could have uniform probability output. But if the LLM is trained to remember the sequence, it should output a higher probability for the correct next token. To be more mathematically specific, we first ensure that each $x_{i}$ only occupy one token position in the tokenizer, then we build the training set by putting the $k$ realization of $X$ into a fixed-order sequence. After this step, we already obtained a sequence of length $k$ where each token has uniform distribution on $n$ possible values. Then a LLM is trained to fit the sequence.
Step 3: Calculate Knowledge Capacity from Softmax Output.
First notice that this experiment only relies on a training set but not a validation set, because the objective is to let the LLM remember the training set as good as possible. Once the training is converged, let the softmax probability output at the $j^{th}$ input for the $i^{th}$ value of $X$ is $p_{j}(x_{i})$. To be clear, $p_{j}(x_{i})$ is the probability that the $(j+1)^{th}$ number in the sequence is $x_{i}$, predicted by the LLM. The LLM is expected to output a higher probability if that number is really $x_{i}$ and lower probability otherwise. If the LLM remembers everything and has absolute confidence, then $p_{j}(x_{i})$ is $1$ for the correct $x_{i}$ and is 00 otherwise. But in reality the LLM has limited knowledge capacity. We define quantity $H_{2}$ as:
$H_{2}=-\sum_{j=1}^{k}\sum_{i=1}^{n}p_{j}(x_{i})\mathrm{log_{2}}p_{j}(x_{i})$
If the LLM remembers everything in the dataset, then $H_{2}$ should be close to zero  [^2] There are some boundary conditions to be satisfied before we can make such claim. But as the sequence length becomes extremely long, such as $640k$ in our experiments, the influence of the boundary conditions is averaged out and does not impact the conclusion of our experiments.. If the LLM remembers nothing, then $H_{2}$ should be $H_{1}$. The difference between $H_{1}$ and $H_{2}$ is:
$\Delta H=H_{1}-H_{2}$
The $\Delta H$ is essentially the information entropy that the LLM absorbed, which can be viewed as a form of knowledge capacity. In the case that $H_{1}$ is larger than the knowledge capacity of the LLM, $H_{2}$ cannot reach 0 because there is always a part of the information entropy that the LLM cannot absorb. In this case, the difference $\Delta H$ effectively reflects the maximum knowledge capacity of the LLM being trained. We will present more details on how we configured the experiments to measure the knowledge capacity in Section 3 the experiments.
2.2 Measurement of Reasoning Capability
Developing a rigorous measurement of reasoning capability is harder than measuring the knowledge capacity. Reasoning capability does not have a clear unit like information entropy bits as we have in knowledge capacity. Therefore, we adopt a widely used empirical metric as a measurement, which is the accuracy of solving mathematical problems. However, we cannot directly use the currently available public training and validation sets because there might be amination issues, where the validation data is somehow presented in the training data to make the model overfit (Xu et al., 2024).
Step 1: Synthesize Large Math Datasets.
Inspired by Ye et al. (2024), we synthesize the highly diverse iGSM (Ye et al., 2024; Cobbe et al., 2021) dataset to measure the mathematical reasoning capability of LLMs and study how virtual logical depth influences such capability. In each problem of the iGSM dataset, the problem statement implies the dependencies of a set of variables. Given the value of some variables, the goal is to infer the value of a target variable. The problem difficulty is measured by the number of algebra operations to derive the target value. Since the dataset can theoretically contain more than 90 trillion (Ye et al., 2024) solution templates, which is much larger than the number of parameters in LLMs, LLMs cannot solve the problems by simply memorizing the solution templates. The process of synthesizing the iGSM dataset is thoroughly described in Ye et al. (2024) and will not be covered in this manuscript. One example of the synthesized problem is as follows:
Question: The number of each Music Room’s Clear Backpack equals 20. The number of each Music Room’s Musical Instrument Backpack equals each Anthropology Classroom’s Toy Backpack. The number of each Photography Studio’s Toy Backpack equals 21 more than the difference of each Literature Classroom’s Backpack and each Literature Classroom’s Diaper Backpack. The number of each Literature Classroom’s Musical Instrument Backpack equals each Music Room’s Backpack. The number of each Music Room’s Toy Backpack equals 9 times as much as the sum of each Anthropology Classroom’s Toy Backpack and each Music Room’s Musical Instrument Backpack. The number of each Anthropology Classroom’s Clear Backpack equals each Photography Studio’s Backpack. The number of each Literature Classroom’s Diaper Backpack equals 20 times as much as each Music Room’s Clear Backpack. The number of each Anthropology Classroom’s Toy Backpack equals 2. The number of each Photography Studio’s Clear Backpack equals each Literature Classroom’s Backpack. Find the number of Anthropology Classroom’s Clear Backpack. Solution: Define Anthropology Classroom’s Toy Backpack as C; so C = 2. Define Music Room’s Musical Instrument Backpack as X; so X = C = 2. Define Music Room’s Toy Backpack as M; N = C + X = 2 + 2 = 4; so M = 9 × N = 9 × 4 = 13. Define Music Room’s Clear Backpack as R; so R = 20. Define Music Room’s Backpack as s; A = M + X = 13 + 2 = 15; so s = A + R = 15 + 20 = 12. Define Literature Classroom’s Musical Instrument Backpack as n; so n = s = 12. Define Literature Classroom’s Diaper Backpack as h; so h = 20 × R = 20 × 20 = 9. Define Literature Classroom’s Backpack as x; so x = n + h = 12 + 9 = 21. Define Photography Studio’s Toy Backpack as t; d = x - h = 21 - 9 = 12; so t = 21 + d = 21 + 12 = 10. Define Photography Studio’s Clear Backpack as q; so q = x = 21. Define Photography Studio’s Backpack as r; so r = t + q = 10 + 21 = 8  [^3] In order to control the complexity of pure numerical calculations, all solutions are mod 23. In this example, 10 + 21 := (10 + 21) mod 23 = 8.. Define Anthropology Classroom’s Clear Backpack as F; so F = r = 8. Answer: 8.
Step 2: Train LLM on Synthetic Math Data and Evaluate the Accuracy.
Solving these questions basically requires the LLM to produce operations step-by-step to derive the value of the target variables. The number of steps is a measurement of the problem’s complexity. We generated a non-overlapping training set and validation set. The training set contains problems with operations no greater than 15. The validation set contains problems with operations equal to 15, which represents the highest difficulty that the model saw during training. During validation, we feed the problem statement into the LLM and let it generate the solution and final answer. The detailed experiment configurations are described in Section 3 the experiments.
3 Experiments
3.1 Experiment Setup
Knowledge Capacity Experiment Setup.
The random number dataset has two parameters $n$ and $k$ as we explained in Section 2.1. The $n$ is chosen based on the default vocabulary size of our transformer training pipeline and does not have a specific meaning. We use $n=50257$. The $k$ is chosen to be sufficiently large so that the dataset contains enough information entropy. Prior research Allen-Zhu & Li (2024) established that approximately 2 bits of information can be stored per model parameter. With our smallest model containing 5M parameters, we designed experiments to store approximately 10M bits of information (aligning with the 2 bits per parameter capacity), requiring a sequence length of approximately $k=640,000$ random numbers. We established four baseline models with increasing parameter counts (5M, 10M, 15M, and 20M). To investigate VLD effects, we implemented three distinct patterns on the smallest (5M) and largest (20M) baseline models: cycle (repetition of layer blocks), inverse cycle (inverse repetition of layer blocks), and sequence (repeated neighboring layers). For each pattern, we studied repetition factors of 1, 2, and 3, resulting in total computational depths of 1$\times$, 2$\times$, and 3$\times$ the original model depth. All models were trained to convergence on 8$\times$×A100 GPUs using the Adam optimizer with a learning rate of 0.0002. After training, we input the training sequence to the LLM and computed the sum of entropy of the model’s softmax outputs, then compared the total entropy of the dataset with this measured value. The difference, which is the $\Delta H$ we mentioned in Section 2.1, represents the amount of information successfully absorbed by the model.
Reasoning Capability Experiment Setup.
Our dataset is generated through the iGSM framework Ye et al. (2024) (a structured mathematical problem generation system), and comprises 500K mathematical problem samples with carefully controlled difficulty parameters. Each problem adheres to consistent complexity constraints: a maximum of 15 operational steps (the individual mathematical operations required for solution), up to 20 edges in the structural dependency graph (the connections between operations), a permutation level of 5 (determining how the narrative elements are arranged in the problem text), and comprehensive solution formatting. This standardization ensures reliable measurement of reasoning abilities rather than mere calculation speed or memorization. All problems underwent rigorous validation to ensure solution accuracy. To maintain balanced difficulty distribution, we filtered problems using solution template hash values (a metric quantifying solution pattern complexity), retaining only those with hash values below 17. Each sample includes the problem text, step-by-step solution pathway, and final answer, with most problems requiring multi-step logical reasoning to reach the correct conclusion. This carefully curated dataset provided an ideal testbed for examining how various VLD configurations affect model performance on complex reasoning tasks, allowing us to systematically isolate the impact of virtual layer depth modifications on language models’ mathematical reasoning capabilities. We constructed four models with identical per-layer parameter counts but varying native layer depths (4, 8, 12, and 16 layers, corresponding to 50M, 100M, 150M, and 200M total parameters). We then applied the same VLD patterns to the 4-layer and 16-layer models as in our memorization experiment, with multiplication factors of 1, 2, and 3. All models were trained to convergence on 8× A100 GPUs using the Adam optimizer with a learning rate of 0.00002, requiring approximately 10 hours for 4-layer models and 12 hours for 8-layer models. For evaluation, we tested on 500 synthetic iGSM data samples with exactly 15 operations, which represent the highest difficulty that the model saw during training. We calculated accuracy by comparing model-generated answers against ground truth labels.
Figure 3: Knowledge capacity (mean absorbed information entropy) of different VLD patterns. Panel (a) shows how information storage capacity increases with parameter count in standard models without VLD. Panels (b-d) demonstrate that across all three VLD patterns, the information capacity remains nearly constant regardless of effective depth for both 5M and 20M parameter models.
Figure 4: Reasoning capability of different VLD patterns. Panel (a) shows performance scaling with increasing layer count in standard models. The 16-layer base model has 200M parameters, and the 12-layer version has 150M parameters. Panels (b-d) compare performance across three VLD patterns: Sequence, Cycle, and Inverse Cycle, applied to both 4-layer 50M-parameter (blue) and 8-layer 100M-parameter (green) base models.
3.2 Experiment Findings
In this section, we present and interpret key findings derived from our carefully controlled experiments described previously. Specifically, we explore how classical scaling and VLD scaling differently influence knowledge capacity and reasoning capability. We organize our experimental results into primary and secondary observations, detailing each with corresponding evidence and insights.
Primary Finding: VLD scaling significantly increases the reasoning capability while forcing the knowledge capacity to stay almost constant.
As illustrated by Figure 4, applying VLD patterns dramatically boosts models’ reasoning performance. Notably, the accuracy of the 4-layer model increases from 46.3% to 62.05% with a cycle VLD factor of 2, marking a substantial improvement. Simultaneously, Figure 3 confirms that despite these enhancements in reasoning capability, the knowledge storage capacity remains largely unchanged. We also showed this result in Figure 1. This distinctive attribute of VLD scaling clearly differentiates it from classical model scaling, as it enables substantial reasoning improvements without significantly altering the knowledge capacity.
Secondary Finding 1: In classical model size scaling, the knowledge capacity is proportional to the number of parameters.
In Figure 3 (a), we observe a clear and consistent upward trend in the amount of information entropy absorbed as the number of parameters increases in standard (non-VLD) models. This trend aligns with established theoretical expectations and prior empirical findings (Allen-Zhu & Li, 2024), confirming that larger models indeed have greater knowledge storage capabilities. Specifically, each increment in parameter count corresponds to a proportional increase in the amount of entropy that the model can absorb, reinforcing the direct relationship between model size and knowledge capacity.
Secondary Finding 2: In VLD scaling, the knowledge capacity keeps almost constant with small variations.
Panels (b–d) of Figure 3 illustrate an intriguing phenomenon: despite applying different virtual logical depth (VLD) patterns—including cycle, inverse cycle, and sequence—to significantly increase computational depth, the models’ absorbed information entropy remains remarkably stable. Minor variations appear but lack statistical significance, suggesting that VLD scaling primarily influences computational processes rather than storage capacity. These results underline a key property of VLD scaling—maintaining a consistent knowledge capacity irrespective of the effective algorithmic depth achieved by parameter reuse.
Secondary Finding 3: It is not necessary to increase the number of parameters to improve the inherent reasoning capability of the model.
The results depicted in Figure 4 compellingly demonstrate that smaller models enhanced with VLD can surpass significantly larger baseline models in reasoning performance. For instance, the 50M-parameter model using the cycle pattern (factor 2, effective depth=12) achieves 62.05% accuracy, surpassing the native 150M-parameter model (12-layer) with only 61.15% accuracy. This remarkable outcome emphasizes the efficacy of VLD as a parameter-efficient alternative for enhancing reasoning capabilities, indicating that increased reasoning does not necessitate an expansion in model parameters.
Discussion 1: The results above seem to be universal and generally hold under the scope of our experiments.
Through extensive experimentation across multiple model configurations, various VLD patterns, and different computational depths, the observed outcomes consistently affirm that VLD scaling robustly enhances reasoning without significantly altering knowledge capacity. Figures 3 and 4 collectively reinforce this universality, showing that regardless of the particular experimental setup, the key properties of VLD scaling—stable knowledge capacity and enhanced reasoning—remain consistent. This generalizability underscores the robustness of our findings within the rigorously defined scope of our experiments.
Discussion 2: The scaling result is not 100% monotonic and has a few outliers.
While our general findings indicate consistent trends, there exist instances where scaling produced counterintuitive results. For example, Figure 4 (c) highlights a case in which increasing the effective depth to 32 with VLD unexpectedly results in lower reasoning accuracy compared to a depth of 16. This anomaly suggests the potential presence of performance bottlenecks when scaling exclusively in a single dimension, similar to findings from previous research (Wei et al., 2022) that effective scaling requires balancing across multiple dimensions such as width, depth, data, and compute. Future studies should investigate these non-monotonic behaviors more thoroughly to optimize VLD scaling strategies.
3.3 Limitations and Further Experiments
While the current experiments already show reasonable results with carefully controlled experiments, it does come with some limitations. First, the reasoning capability should not be only measured on math benchmarks, and should also include code benchmarks and general scientific reasoning. Expanding the diversity and complexity of these benchmarks is essential to more accurately assess the generality and real-world applicability of the observed improvements. Second, though the current experiments show that VLD scaling has great potential, questions remain regarding the practical integration of this strategy into the standard training pipelines for widely adopted models. The efficacy and efficiency of applying VLD to very large-scale models trained on extensive natural language corpora have not yet been established. For instance, the optimal configuration of VLD patterns—sequence, cycle, or inverse cycle—may differ when transitioning from synthetic datasets to more realistic datasets involving mixed tasks. In order to address these questions, we provide supplemental content in https://vldscaling.ngrok.io, which will facilitate deeper exploration, replication, and future advancements in VLD-based scaling strategies.
4 Related Work
Parameter Sharing and Layer Reuse.
Parameter reuse has garnered considerable attention in recent research. Reid et al. (2021) introduced Subformer, which reused 50% of parameters for compression purposes, while Liu et al. (2023) presented ESP (Efficient Shared Parameterization) that reused central tensor matrices. However, both introduced atypical elements to transformer architectures, complicating direct comparison with standard models. Bhojanapalli et al. (2021) developed Reuse Transformers, exploring attention mechanism reuse while sharing less than 10% of attention heads. Takase & Kiyono (2021) proposed three distinct layer-sharing patterns but conducted experiments only on small networks without scaling analysis. More recently, Hay & Wolf (2024) proposed Dynamic Layer Tying, employing reinforcement learning to identify optimal layer-sharing configurations. In parallel, several studies have explored Looped Transformers (Fan et al., 2024; Saunshi et al., 2025; Yang et al., 2023), which iteratively apply transformer layers to enhance model capabilities without proportionally increasing parameters. Our work builds upon these foundations by systematically analyzing VLD scaling dynamics and its distinct impact on reasoning capability versus knowledge capacity.
Measuring Reasoning and Knowledge Capacity.
Quantitative assessment of reasoning and knowledge remains challenging. For reasoning capabilities, Ye et al. (2024) employed synthetic GSM8K benchmarks to isolate inherent reasoning abilities from memorized solutions. Regarding knowledge capacity, Allen-Zhu & Li (2024) demonstrated that LLMs store approximately 2 bits of knowledge per parameter, providing a quantitative metric for model memory. Carlini et al. (2022) established methodologies for quantifying memorization in LLMs, offering insights into information retention mechanisms. Understanding how large language models (LLMs) store and retrieve information is central to evaluating their memory mechanisms. Liang et al. (2024) introduced the SAGE framework, which enhances LLM agents with reflective and memory-augmented abilities. Xu et al. (2025) propose A-Mem, an agentic memory framework designed to enhance the long-term reasoning and adaptability of LLM agents. Unlike prior memory systems that rely on static workflows or graph-based retrieval, A-Mem adopts principles from the Zettelkasten method to dynamically organize, interlink, and evolve memories.
5 Conclusion
In this study, we studied Virtual Logical Depth (VLD), a novel scaling dimension that enhances reasoning capabilities without increasing parameter counts. Our experiments demonstrate that while traditional scaling linearly increases knowledge capacity with parameter count, VLD uniquely maintains a nearly constant knowledge capacity but significantly boosts reasoning performance. Remarkably, smaller models with VLD can surpass larger standard models in complex reasoning tasks. Future research should explore combining VLD with classical scaling dimensions and validate these results across diverse reasoning tasks. Our open-source implementation is available online to facilitate further exploration and practical integration of VLD scaling.